OpenStudy (anonymous):

Solve: x/2 + 3/2 = 1 a. x = -2 b. x = -1 c. x = 1 d. x = 2

OpenStudy (lalaly):

x+3=2 x=-1 so b

OpenStudy (anonymous):

(x+3)/2 = 1 Multiply both sides by 2. Now you have: x+3 = 2 x = -1 by basic algebra. So your answer is b. Hope this helps. If it does, please choose good answer. Visit http://www.tutorsean.net for free math videos.